b. 1.2.3, 4 only analysis the NWC for Discontinued Operations sums years: c. 0,1,2,3,4, 5 only d. 0,1, 2, 3, 4 only d bevelop answer all for thumbs up
Answer #1

27: d

The NWC begins with year 0 investment and ends with recovery at the end of the project which is year 4 in this case.

28:a

All cash flows are to be accounted for while doing the NPV analysis.

29:Option 2

Change in year 0 = 0

Year 1 = 100-50= $50

Year 2 = 120-50 = 70

Year 3 = 150-50 = 100